Assume that the world consists of 2 countries ,Ghana and Nigeria and these countries do not read with each other and there are only 2 goods In this world-Gold and Crude oil.Assume that Ghana needs 12 workers to produce a unit of Gold and 6 workers to produce a barrel of Crude oil.Nigeria needs 2 workers to produce unit of Gold and 4 workers to produce a barrel of Crude oil. Assume that both countries have 48 workers each.Usingbthe Ricardian model,answer this.
Assume the countries are now allowed to trade with each other.Assume that the equilibrium with trade is such that one country fully specializes in the production of Gold and the other country in Crude oil,Which country specializes in which good?
1
Expert's answer
2020-10-21T10:18:09-0400

In this case Ghana can produce either 48/12 = 4 units of gold or 48/6 = 8 barrels of oil, Nigeria can produce either 48/2 = 24 units of gold or 48/4 = 12 barrels of oil.

So, Nigeria will produce gold, and Ghana will produce oil, as they have comparative advantage in producing these goods.